question

"Consider the following data on an asset
Cost of the asset, I = $87,000
Useful life, N = 7 years
What is the book value at the end of the useful life if you depreciate according to the double-declining-balance (DDB) method?"
Answer #1

Straight line depreciation rate= 1/7= 14.28%

Double declining balance method depreciation rate= 2*14.28= 28.6%

Year Book value at the beginning of year Depreciation rate Depreciation expense Book value at the end of year
1 $87,000 28.6% $24,882 $62,118
2 $62,118 28.6% $17,765.75 $44,352.25
3 $44,352.25 28.6% $12,684.74 $31,667.51
4 $31,667.51 28.6% $9,056.91 $22,610.6
5 $22,610.6 28.6% $6,466.63 $16,143.97
6 $16,143.97 28.6% $4,617.18 $11,526.79
7 $11,526.79 $11,526.79 $0